# Reflexive homology and involutive Hochschild homology as equivariant Loday constructions

### v3

@@ 6. Relating $\mathcal{L}_{S^{\sigma}}^{C_{2}}({\underline{R}})$ to reflexive homology
With this choice of sign, the map $r$ is a chain map, so the quotient by $1-r$ is still a chain complex. In the following we denote by $B^{k}_{*}(A,A\otimes A^{\mathrm{op}},M)$ the chain complex associated to the simplicial $k$-module $B^{k}(A,A\otimes A^{\mathrm{op}},M)$.
@@
The second isomorphism follows from the fact that $2$ is invertible in both complexes: As $2$ is invertible in $A$, the unit of $A$, $k\rightarrow A$ factors through $k[\frac{1}{2}]$. We can express every level of each of the complexes as the direct sum of the $+1$-eigenspace and the $-1$-eigenspace of the action of the generator of $C_{2}$ on them. Since the actions commute with $d$, in fact each of the complexes breaks up as the direct sum of a positive subcomplex and a negative subcomplex.
@@
This is a simplicial $A\otimes A^{\mathrm{op}}$-module map, and covers the identity on the $A$ being resolved since in level 00 it sends $a_{0}\otimes(a_{1}\otimes a_{2})$ to $a_{2}a_{0}\otimes a_{1}$ and both of these map down to $a_{2}a_{0}a_{1}\in A$. This map is not $C_{2}$-equivariant, but if we define $g:=r\circ f\circ\tau$, we get $g_{n}\colon B^{k}_{n}(A,A\otimes A^{\mathrm{op}},A\otimes A^{\mathrm{op}})\to B^{k}_{n}(A,A,A)$ with
@@
This is not a simplicial map but it is an $A\otimes A^{\mathrm{op}}$-module map and it is a chain map since $r$, $f$, and $\tau$ are chain maps. Again, it covers the identity on $A$ since on level 00, $a_{0}\otimes(a_{1}\otimes a_{2})\mapsto a_{2}\otimes a_{0}a_{1}$ and both of these map down to $a_{2}a_{0}a_{1}\in A$.
@@
Since $\phi$ is a quasi-isomorphism, its mapping cone, $\mathrm{cone}(\phi)$, is acyclic. The mapping cone is also a bounded-below chain complex of flat right $R$-modules, so it can be viewed as a flat resolution of the 00-module, possibly with a shift. We suspend it, so that $\Sigma^{a}\mathrm{cone}(\phi)$ is a non-negative chain complex whose bottom chain group is in degree zero. Since flat resolutions can be used to calculate $\mathsf{Tor}$,
@@
Taking our identification of $\mathcal{L}_{S^{\sigma}}^{C_{2}}({\underline{R}})$ with ${\underline{B(R,R\otimes R,R)}}^{\mathrm{fix}}$ from Theorem 5.1 together with Corollary 6.2 we obtain the following comparison result between the homology groups of the $C_{2}$-Loday construction for the flip-circle $S^{\sigma}$ and ${\underline{R}}^{\mathrm{fix}}$ on the one hand and the reflexive homology groups on the other hand:
@@ 7. Involutive Hochschild homology as a Loday construction
[…]
@@
On the other hand, if $G$ is a group with $n$ elements $\{g_{1},\ldots,g_{n}\}$, we can consider the unreduced suspension of $G$, $SG$. This is the graph
@@
8. The cases ${\underline{\mathbb{F}_{2}}}^{c}$ and ¯${\underline{\mathbb{Z}}}^{c}$
@@ 8.1. Comparison for ${\underline{\mathbb{F}_{2}}}^{c}$
$[1\otimes 1\otimes 1]=[\mathsf{res}(1)\otimes 1\otimes 1]\sim 1\otimes\mathsf{tr}(1\otimes 1)=2\cdot 1\otimes 1\otimes 1=0$
@@
8.2. Comparison for ¯${\underline{\mathbb{Z}}}^{c}$
We consider the ring of integers and this only carries a trivial $C_{2}$-action. We know that norm restriction of ¯${\underline{\mathbb{Z}}}^{c}$ gives the $C_{2}$-Burnside Tambara functor, $N_{e}^{C_{2}}i_{e}^{*}{\underline{\mathbb{Z}}}^{c}\cong{\underline{A}}$. This is the monoidal unit for the $\Box$-product. We showed in [lrz, Lemma 5.1] that for two arbitrary commutative rings $A$ and $B$, ${\underline{A}}^{c}\Box{\underline{B}}^{c}\cong{\underline{(A\otimes B)}}^{c}$ and hence
¯${\underline{\mathbb{Z}}}^{c}\Box{\underline{\mathbb{Z}}}^{c}\cong{\underline{(\mathbb{Z}\otimes\mathbb{Z})}}^{c}\cong{\underline{\mathbb{Z}}}^{c}.$
@@ 8.2. Comparison for ¯${\underline{\mathbb{Z}}}^{c}$
where the right-hand side denotes the constant simplicial $C_{2}$-Tambara functor with value ¯${\underline{\mathbb{Z}}}^{c}$.
@@
$\mathcal{L}_{S^{\sigma}}^{C_{2}}({\underline{\mathbb{Z}}}^{c})_{n}$ =¯$={\underline{\mathbb{Z}}}^{c}\Box(N_{e}^{C_{2}}i_{e}^{*}{\underline{\mathbb{Z}}}^{c})^{\Box n}\Box{\underline{\mathbb{Z}}}^{c}$
¯$\cong{\underline{\mathbb{Z}}}^{c}\Box{\underline{\mathbb{Z}}}^{c}$
¯$\cong{\underline{\mathbb{Z}}}^{c}.$
@@
For the $C_{2}$-Tambara functor ¯${\underline{\mathbb{Z}}}^{c}$ the homotopy groups
@@
We saw above that $\pi_{*}(\mathcal{L}_{S^{\sigma}}^{C_{2}}({\underline{\mathbb{Z}}}^{c}))(C_{2}/C_{2})$ is concentrated in degree =0*=0∗ = 0 with value $\mathbb{Z}$ whereas $\mathsf{HR}^{+,\mathbb{Z}}_{*}(\mathbb{Z})$ and $\mathsf{iHH}^{\mathbb{Z}}_{*}(\mathbb{Z})$ both give $H_{*}(C_{2};\mathbb{Z})$. ∎
@@ 9. The case of rings and algebras with anti-involution
We assume now that $R$ is an associative ring with anti-involution. In this case the fixed point Mackey functor ${\underline{R}}^{\mathrm{fix}}$ is not an associative Green functor: For $a,b\in R^{C_{2}}$ we get that $\tau(ab)=\bar{b}\bar{a}=ba$ and as $a$ and $b$ do not necessarily commute, $ab$ is not, in general, a fixed point. But $\underline{R}^{\mathrm{fix}}$ does carry the structure of a discrete $E_{\sigma}$-ring [akgh, Example 6.12] and this in turn is a Hermitian Mackey functor in the sense of [do, Definition 1.1.1]
@@
$(\tilde{N}^{C_{2}}_{e}i^{*}_{e}\underline{R}^{\mathrm{fix}})\Box\underline{R}^{\mathrm{fix}}=\begin{cases}\bigl{(}(\mathbb{Z}\{R\}\oplus(R\otimes R^{\mathrm{op}})/C_{2})/\mathrm{TR}\otimes R^{C_{2}}\oplus(R\otimes R^{\mathrm{op}}\otimes R)/C_{2}\bigr{)}/\mathrm{FR}&\text{ at }C_{2}/C_{2}\\ R\otimes R^{\mathrm{op}}\otimes R&\text{ at }C_{2}/e.\end{cases}$
@@
    $[a\otimes b]\otimes x\mapsto axb+\bar{b}x\bar{a}.$
@@
    The $C_{2}$-action on $R\otimes R^{\mathrm{op}}\otimes R$ sends a generator $a\otimes b\otimes y$ to $\bar{b}\otimes\bar{a}\otimes\bar{y}$. We send a $C_{2}$-equivalence class $[a\otimes b\otimes y]$ to $ayb+\bar{b}\bar{y}\bar{a}$.
@@
A direct inspection shows that $[a\otimes b]\otimes x$ and $[\bar{b}\otimes\bar{a}]\otimes x$ map to the same element. Similarly, the value on $[a\otimes b\otimes y]$ and $[\bar{b}\otimes\bar{a}\otimes\bar{y}]$ agrees. It is also straightforward to see that the module structure respects Tambara reciprocity. For Frobenius reciprocity we have to compare three expressions:
  • $[a\otimes b]\otimes x$ is $\mathsf{tr}(a\otimes b)\otimes x$ and this is identified with $[a\otimes b\otimes\mathsf{res}(x)]=[a\otimes b\otimes x]$. Both terms are mapped to $axb+\bar{b}x\bar{a}$ because $x$ is a fixed point.
  • A term $\{a\}\otimes\mathsf{tr}(y)$ is sent to $a(y+\bar{y})\bar{a}$. It is identified with $[\mathsf{res}\{a\}\otimes y]=[a\otimes\bar{a}\otimes y]$ and this goes to $ay\bar{a}+a\bar{y}\bar{a}$.
@@
    $[a\otimes b]\otimes(y+\bar{y})=[a\otimes b]\otimes\mathsf{tr}(y)=\mathsf{res}([a\otimes b])\otimes y.$
@@
to $\underline{R}^{\mathrm{fix}}$. At the free level we send $a\otimes(b\otimes c)$ to $cab$ and this propagates to the trivial level where we map $[y\otimes a\otimes b]$ to $bya+\bar{a}\bar{y}\bar{b}$ and $x\otimes[a\otimes b]$ to $bxa+\bar{a}x\bar{b}$. A term $x\otimes\{a\}$ goes to $\bar{a}xa$. Then a proof dual to the above shows that this indeed gives a well-defined right module structure and that this right-module structure is compatible with the left-module structure so that we actually obtain a bimodule structure. ∎
@@
With this we can define $\mathcal{L}_{S^{\sigma}}^{C_{2},{\underline{k}}^{c}}({\underline{A}}^{\mathrm{fix}})$ for an associative $k$-algebra $A$ with anti-involution and obtain an isomorphism of simplicial $C_{2}$-Mackey functors
@@
We have to adapt the statement and the proof of Proposition 4.1 and claim that for an associative $k$-algebra $A$ with anti-involution we obtain that

### v4

@@ 6. Relating $\mathcal{L}_{S^{\sigma}}^{C_{2}}({\underline{R}})$ to reflexive homology
With this choice of sign, the map $r$ is a chain map, so the quotient by 1r1-r is still a chain complex. In the following we denote by $B^{k}_{*}(A,A\otimes A^{\mathrm{op}},M)$ the chain complex associated to the simplicial $k$-module $B^{k}(A,A\otimes A^{\mathrm{op}},M)$.
@@
The second isomorphism follows from the fact that $2$ is invertible in both complexes: As $2$ is invertible in $A$, the unit of $A$, $k\rightarrow A$ factors through $k[\frac{1}{2}]$. We can express every level of each of the complexes as the direct sum of the +1+1-eigenspace and the 1-1-eigenspace of the action of the generator of $C_{2}$ on them. Since the actions commute with $d$, in fact each of the complexes breaks up as the direct sum of a positive subcomplex and a negative subcomplex.
@@
This is a simplicial $A\otimes A^{\mathrm{op}}$-module map, and covers the identity on the $A$ being resolved since in level 0 it sends $a_{0}\otimes(a_{1}\otimes a_{2})$ to $a_{2}a_{0}\otimes a_{1}$ and both of these map down to $a_{2}a_{0}a_{1}\in A$. This map is not $C_{2}$-equivariant, but if we define $g:=r\circ f\circ\tau$, we get $g_{n}\colon B^{k}_{n}(A,A\otimes A^{\mathrm{op}},A\otimes A^{\mathrm{op}})\to B^{k}_{n}(A,A,A)$ with
@@
This is not a simplicial map but it is an $A\otimes A^{\mathrm{op}}$-module map and it is a chain map since $r$, $f$, and $\tau$ are chain maps. Again, it covers the identity on $A$ since on level 0, $a_{0}\otimes(a_{1}\otimes a_{2})\mapsto a_{2}\otimes a_{0}a_{1}$ and both of these map down to $a_{2}a_{0}a_{1}\in A$.
@@
Since $\phi$ is a quasi-isomorphism, its mapping cone, $\mathrm{cone}(\phi)$, is acyclic. The mapping cone is also a bounded-below chain complex of flat right $R$-modules, so it can be viewed as a flat resolution of the 0-module, possibly with a shift. We suspend it, so that $\Sigma^{a}\mathrm{cone}(\phi)$ is a non-negative chain complex whose bottom chain group is in degree zero. Since flat resolutions can be used to calculate $\mathsf{Tor}$,
@@
Taking our identification of $\mathcal{L}_{S^{\sigma}}^{C_{2}}({\underline{R}})$ with ${\underline{B(R,R\otimes R,R)}}^{\mathrm{fix}}$ from Theorem 5.1 together with Corollary 6.2 we obtain the following comparison result between the homology groups of the $C_{2}$-Loday construction for the flip circle $S^{\sigma}$ and ${\underline{R}}^{\mathrm{fix}}$ on the one hand and the reflexive homology groups on the other hand:
@@ 7. Involutive Hochschild homology as a Loday construction
$\ldots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${\pi_{n}BM_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{n}BM_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{n}BM_{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{n-1}BM_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\ldots}$
@@
On the other hand, if $G$ is a group with $n$ elements $\{g_{1},\ldots,g_{n}\}$, we can consider the unreduced suspension of $G$, SGSG. This is the graph
@@
8. The cases ${\underline{\mathbb{F}_{2}}}^{c}$ and ${\underline{\mathbb{Z}}}^{c}$
@@ 8.1. Comparison for ${\underline{\mathbb{F}_{2}}}^{c}$
$1\otimes 1\otimes 1]=[\mathsf{res}(1)\otimes 1\otimes 1]\sim 1\otimes\mathsf{tr}(1\otimes 1)=2\cdot 1\otimes 1\otimes 1=0$
@@
8.2. Comparison for ${\underline{\mathbb{Z}}}^{c}$
We consider the ring of integers and this only carries a trivial $C_{2}$-action. We know that norm restriction of ${\underline{\mathbb{Z}}}^{c}$ gives the $C_{2}$-Burnside Tambara functor, $N_{e}^{C_{2}}i_{e}^{*}{\underline{\mathbb{Z}}}^{c}\cong{\underline{A}}$. This is the monoidal unit for the $\Box$-product. We showed in [lrz, Lemma 5.1] that for two arbitrary commutative rings $A$ and $B$, ${\underline{A}}^{c}\Box{\underline{B}}^{c}\cong{\underline{(A\otimes B)}}^{c}$ and hence
${\underline{\mathbb{Z}}}^{c}\Box{\underline{\mathbb{Z}}}^{c}\cong{\underline{(\mathbb{Z}\otimes\mathbb{Z})}}^{c}\cong{\underline{\mathbb{Z}}}^{c}.$
@@ 8.2. Comparison for ${\underline{\mathbb{Z}}}^{c}$
where the right-hand side denotes the constant simplicial $C_{2}$-Tambara functor with value ${\underline{\mathbb{Z}}}^{c}$.
@@
$\mathcal{L}_{S^{\sigma}}^{C_{2}}({\underline{\mathbb{Z}}}^{c})_{n}$ $={\underline{\mathbb{Z}}}^{c}\Box(N_{e}^{C_{2}}i_{e}^{*}{\underline{\mathbb{Z}}}^{c})^{\Box n}\Box{\underline{\mathbb{Z}}}^{c}$
$\cong{\underline{\mathbb{Z}}}^{c}\Box{\underline{\mathbb{Z}}}^{c}$
$\cong{\underline{\mathbb{Z}}}^{c}.$
@@
For the $C_{2}$-Tambara functor ${\underline{\mathbb{Z}}}^{c}$ the homotopy groups
@@
We saw above that $\pi_{*}(\mathcal{L}_{S^{\sigma}}^{C_{2}}({\underline{\mathbb{Z}}}^{c}))(C_{2}/C_{2})$ is concentrated in degree =0*=0 with value $\mathbb{Z}$ whereas $\mathsf{HR}^{+,\mathbb{Z}}_{*}(\mathbb{Z})$ and $\mathsf{iHH}^{\mathbb{Z}}_{*}(\mathbb{Z})$ both give $H_{*}(C_{2};\mathbb{Z})$. ∎
@@ 9. The case of rings and algebras with anti-involution
We assume now that $R$ is an associative ring with anti-involution. In this case the fixed point Mackey functor ${\underline{R}}^{\mathrm{fix}}$ is not an associative Green functor: For $a,b\in R^{C_{2}}$ we get that $\tau(ab)=\bar{b}\bar{a}=ba$ and as $a$ and $b$ do not necessarily commute, $ab$ is not, in general, a fixed point. But $\underline{R}^{\mathrm{fix}}$ does carry the structure of a discrete $E_{\sigma}$-ring [akgh, Example 6.12] and it is also a Hermitian Mackey functor in the sense of [do, Definition 1.1.1]
@@
[…]
@@
    $a\otimes b]\otimes x\mapsto axb+\bar{b}x\bar{a}.$
@@
    The $C_{2}$-action on $R\otimes R^{\mathrm{op}}\otimes R$ sends a generator $a\otimes b\otimes y$ to $\bar{b}\otimes\bar{a}\otimes\bar{y}$. We send a $C_{2}$-equivalence class $a\otimes b\otimes y]$ to $ayb+\bar{b}\bar{y}\bar{a}$.
@@
A direct inspection shows that $a\otimes b]\otimes x$ and $\bar{b}\otimes\bar{a}]\otimes x$ map to the same element. Similarly, the value on $a\otimes b\otimes y]$ and $\bar{b}\otimes\bar{a}\otimes\bar{y}]$ agrees. It is also straightforward to see that the module structure respects Tambara reciprocity. For Frobenius reciprocity we have to compare three expressions:
  • $a\otimes b]\otimes x$ is $\mathsf{tr}(a\otimes b)\otimes x$ and this is identified with $a\otimes b\otimes\mathsf{res}(x)]=[a\otimes b\otimes x]$. Both terms are mapped to $axb+\bar{b}x\bar{a}$ because $x$ is a fixed point.
  • A term $\{a\}\otimes\mathsf{tr}(y)$ is sent to $a(y+\bar{y})\bar{a}$. It is identified with $\mathsf{res}\{a\}\otimes y]=[a\otimes\bar{a}\otimes y]$ and this goes to $ay\bar{a}+a\bar{y}\bar{a}$.
@@
    $a\otimes b]\otimes(y+\bar{y})=[a\otimes b]\otimes\mathsf{tr}(y)=\mathsf{res}([a\otimes b])\otimes y.$
@@
to $\underline{R}^{\mathrm{fix}}$. At the free level we send $a\otimes(b\otimes c)$ to cabcab and this propagates to the trivial level where we map $y\otimes a\otimes b]$ to $bya+\bar{a}\bar{y}\bar{b}$ and $x\otimes[a\otimes b]$ to $bxa+\bar{a}x\bar{b}$. A term $x\otimes\{a\}$ goes to $\bar{a}xa$. Then a proof dual to the above shows that this indeed gives a well-defined right module structure and that this right-module structure is compatible with the left-module structure so that we actually obtain a bimodule structure. ∎
@@
With this we can define $\mathcal{L}_{S^{\sigma}}^{C_{2},{\underline{k}}^{c}}({\underline{A}}^{\mathrm{fix}})$ for an associative $k$-algebra AA with anti-involution and obtain an isomorphism of simplicial $C_{2}$-Mackey functors
@@
We have to adapt the statement and the proof of Proposition 4.1 and claim that for an associative $k$-algebra AA with anti-involution we obtain that